# DriveSOTIF: Advancing SOTIF Through Multimodal Large Language Models

### v2

@@ IV-C2 Quantitative Analysis
Fine-tuned models show consistent and significant improvements across standard metrics. In image captioning, the fine-tuned Blip2 6.7B model achieves ROUGE-L, CIDEr, and SPICE scores of 0.421, 0.536, and 0.155, representing increases of 24.6%, 204.5%, and 61.5% over the baseline, respectively. Similar improvements are observed across BLEU, METEOR, and other benchmarks, indicating enhanced contextual relevance and descriptive accuracy. Notably, the smaller Blip2 2.7B model also performs competitively, highlighting the value of balancing model size and efficiency.
For VQA, larger models generally outperform smaller ones, with the 8B variant of InternVL3 reaching the highest baseline accuracy at 59.1%. Fine-tuning yields substantial accuracy gains for all models: for instance, InternVL3 2B rises from 58.0% to 64.8%. Trustworthiness, clarity, and coherence metrics—scored by LLM-as-judge—also improve, often by more than 0.6 points on a 5-point scale. Of note, fine-tuned Qwen2-VL 2B achieves competitive performance (61.01% accuracy) with only a fraction of the parameters compared to larger models. While accuracy and most quality metrics improve substantially, some metrics such as “relevance” show smaller gains, that even baseline models can identify broad risk factors. However, they lack the detailed, domain-specific knowledge required for precise safety recommendations.
Both qualitative and quantitative results confirm that supervised fine-tuning on the DriveSOTIF dataset significantly enhances model performance for perception-related SOTIF scenarios. The domain-specific dataset enables models to better interpret complex, safety-critical driving situations, supporting more robust and reliable reasoning for autonomous driving applications.
@@ V-A Model Deployment and Real-time Performance Evaluation
MLLMs can be deployed across various environments, including local devices, cloud platforms, and edge computing systems such as multi-access edge computing (MEC). Common deployment frameworks include Huggingface, TensorRT-based solutions [43], and the vLLM framework [44], which was selected for this research due to its efficiency and support for advanced scheduling strategies.
@@
Two inference strategies were considered. Sequential inference processes inputs one at a time as they arrive. This provides deterministic, per-sample latency—an essential requirement for real-time, safety-critical applications like autonomous driving. Continuous batching, by contrast, dynamically aggregates requests and schedules token-level processing across partially completed sequences. Instead of traditional static batching, vLLM uses iteration-level scheduling and PagedAttention to refill GPU slots as soon as they are available. This maximizes GPU utilization and minimizes token-level idle time, resulting in higher throughput without sacrificing latency [44].
@@ V-A1 Sequential Processing
We tested this mode using 111 images from DriveSOTIF, each paired with a prompt for SOTIF risk analysis. Generation time for each image were recorded and aggregated in Table V-A1. GPU VRAM usage was measured via HuggingFace APIs due to vLLM’s memory pre-allocation behavior.
[b]s
@@
    LMDeploy is used for InternVL3 deployment.
The peak GPU memory usage and processing time per image are recorded. The statistical analysis is performed, and the result is presented in Table V-A1. Based on our results, quantization reduces GPU memory usage while improving real-time performance. Specifically, Qwen2-VL (2B) in 4-bit quantization achieves the lowest memory footprint (2.05 GB) but with a slight latency increase compared to FP8. FP8 quantization yields the best speed on Jetson Orin, reducing inference time by approximately 21% relative to the full-precision model. For larger models like Qwen2.5-VL (3B), 4-bit quantization significantly reduces memory usage from 7.64 GB to 2.90 GB and improves inference speed on both hardware platforms, achieving over 1.3× speedup on Jetson Orin.
Comparing across models, InternVL3-1B consistently exhibits strong performance with the lowest latency and compact memory usage, making it suitable for deployment in resource-constrained environments. However, as the model size increases (e.g., InternVL3-2B), the latency grows accordingly, especially on edge devices.
@@ V-A2 Continuous Processing for high-speed driving
To evaluate continuous inference, we replayed four 30 Hz video streams of high-speed driving scenarios, triggering model inference on every frame. Each frame generates a request to the InternVL3-1B vision-language model, which is deployed on an RTX 3090 with continuous batching enabled  [^4] Please refer to Section V of the supplemental material for results on RTX4090 and duel RTX4090 setup. The system supports multiple requests in parallel, and both queuing and auto-timeout mechanisms are in place: when more requests arrive than the hardware can process immediately, excess requests are placed in a first-in, first-out (FIFO) queue, and any request waiting or processing beyond a set timeout threshold is automatically dropped. We varied the maximum number of in-flight (active) requests—from 1 up to 30 (concurrency)—and, for each setting, recorded the API response times and the number of active requests throughout the experiment (see Figure 3).
@@
We visualized these dynamics by jointly plotting queue wait times and API response times for each concurrency setting (Figure 4). At low concurrency, the system processes requests quickly once they reach the model, but incoming requests rapidly accumulate in the queue, leading to prolonged wait times and, frequently, discarded requests due to timeouts. As concurrency increases, response times become more variable and overall latency rises, but queue wait times decrease since more requests are processed in parallel. In contrast, with very low concurrency,
This illustrates a fundamental tradeoff between response time and queue waiting time with various concurrency levels. High concurrency leads to longer per-request processing times but keeps queue wait times low. Low concurrency minimizes processing latency for individual requests but causes significant queuing and dropped requests under heavy load. For real-time applications such as autonomous driving, it is essential to maintain concurrency at a moderate level to balance response latency and throughput, ensuring timely and reliable perception without excessive delays or data loss.
@@
Figure 4: Tradeoff between API reponse time and queue waiting time in continuous inference mode.
@@ V-C Results
As shown in Table V-B, in our experiments, fine-tuned MLLMs performed effectively in challenging environments, yet it also highlighted critical areas for enhancement. Notably, in Waterloo’s snowy conditions, the MLLM successfully managed complex scenarios with obscured road markers and temporary obstacles like snowplows and maintenance crews. This performance highlights the model’s robust feature recognition capabilities under adverse weather conditions, proving its potential in dynamic real-world applications.
@@
Conversely, the MLLM’s performance in detecting subtle hazards, such as the partially visible child in a box during the China experiments, revealed some limitations. While the model detected unusual patterns associated with the obscured visual cues, it did not accurately capture and interpret the partially or fully obscured child. This highlights a scenario that is challenging even for human drivers. One possible reason is the current capacity of the visual encoders used in the MLLM.
@@ VI-A Implications of Experimental Results
The experimental findings highlight three key capabilities: accurate risk detection and cognition under adverse condition, causal reasoning linking environmental factors to safety outcomes, and actionable mitigation strategies that address specific SOTIF vulnerabilities. With only 2 GB of GPU memory, the fine-tuned InternVL3-1B model can be deployed to vehicles and replace the API-based vision model for local (near) real-time visual risk assessment. This eliminates privacy concerns, cybersecurity risks, and reliance on internet access.
Pairing this model with a decision-making framework like DriveLLM[45] enables autonomous systems to incorporate SOTIF risks into real-time decision-making. This integration improves safety in complex environments and potentially prevents fatal accidents like the child-hidden-in-the-box scenario, where conventional perception systems may fail to anticipate hidden risks.
Most importantly, deploying a fine-tuned model that supports continuous improvement significantly enhances adaptability. By continuously updating the dataset and applying techniques like SFT and reinforcement learning from human feedback (RLHF) [47], the MLLM can progressively refine its performance and extend to other tasks such as scenario understanding and ODD assessment.
@@ VI-B Domain Transfer
From the labeled response presented in Table III, proprietary LLM models demonstrate the ability to provide accurate and relevant answers to complex, domain-specific questions related to AD perception and SOTIF risks, even without prior specialized training. This highlights their effectiveness in zero-shot and one-shot scenarios, where little or no task-specific data is available.
@@ VI-C Hallucination
To address this challenge, we implemented a multi-stage hallucination suppression strategy during the SFT process. Our approach includes: (1) a post-processing validation step using an independent LLM to cross-verify generated responses against input images, (2) consistency filtering that removes outputs failing validation checks, and (3) iterative correction of inconsistent annotations. This targeted intervention significantly reduces hallucinated content entering the fine-tuning pipeline, ensuring our dataset maintains grounding in factual, image-based annotations.
However, inference-time hallucination remains a persistent challenge, particularly due to inherent knowledge gaps and biases in pre-trained model backbones. Existing mitigation strategies present distinct trade-offs for real-time applications. Retrieval-augmented generation [49], while effective at grounding responses in factual content, introduces additional inference latency that may be incompatible with the sub-second response requirements of autonomous driving systems. Human feedback mechanisms [50], though widely adopted in industry and demonstrably effective at improving model reliability, face scalability constraints when applied to high-frequency inference scenarios typical in autonomous vehicles. At the moment, one feasible direction for addressing inference-time hallucination is grounding model response through a human-in-loop semi-auto feedback mechanism. Other methods, like latent space steering [51], can also be explored to reduce hallucination while maintaining real-time performance.
@@ VI-D Future of LLM in SOTIF
As current research on LLMs in SOTIF applications remains limited, this section outlines potential future uses of both text-based and multimodal LLMs. First, a universal description system for SOTIF scenarios can be developed to provide standardized descriptions. Building on this foundation, custom-trained LLMs can generate JSON-based representations for scenario detection, communication, and risk identification. This standardization would make it easier to share and compare SOTIF scenarios across different research groups and industry applications.
@@ VII Conclusion and Future Work
Our research demonstrates that SFT techniques applied to MLLMs using the DriveSOTIF dataset can enhance perception-related SOTIF risk detection in autonomous driving systems. By introducing a small domain-specific dataset intended for model fine-tuning, we were able to evaluate strengths and limitations of MLLMs in perception-related SOTIF problems. Benchmark results indicate that our fine-tuned MLLMs outperform baseline models both qualitatively and quantitatively.
These improvements enhance risk identification and situational awareness, even under challenging conditions such as adverse weather and occlusion-heavy scenarios. This is confirmed by experiments on data collected from both China and Canada. Real-world performance testing further confirms that the deployed models achieve an average inference time of 0.59 second per image, making them suitable for low- to moderate-speed applications like urban navigation, automated parking, and traffic jam assistance.
More importantly, by integrating fine-tuned MLLMs with LLM-based decision-making frameworks, autonomous vehicles can incorporate perception risks into real-time operations, ultimately improving road safety. Moreover, the ability of MLLMs to interpret complex scenarios through natural language reasoning establishes concrete foundations for transparent safety assessment through natural language explanations of risk factors and mitigation strategies. Lastly, the on-device and MEC-based deployment of these lightweight models can lead to a privacy-preserving, onboard-cloud hybrid LLM deployment for next-generation autonomous vehicles.
For future work, we plan to improve real-time performance for high-speed safety-critical applications, and minimize hallucinations during inference stage. This will require interdisciplinary collaboration spanning natural language processing and autonomous driving to realize the full potential of MLLMs in safer autonomous driving systems.

### v3

@@ IV-C2 Quantitative Analysis
Fine-tuning improves performance across all evaluation metrics. For image captioning, the fine-tuned BLIP-2 6.7B model obtained the best results, achieving ROUGE-L, CIDEr, and SPICE scores of 0.421, 0.536, and 0.155, respectively. This marks an increase of 24.56%, 204.54%, and 61.46% compared to the baseline version. Similar improvements were observed in BLEU-4 and METEOR scores, indicating models can generate more accurate and contextually relevant descriptions of complex scenes. Notably, the fine-tuned BLIP-2 2.7B model achieved comparable performance.
For the VQA task, larger models generally outperform smaller ones. The InternVL3-8B model achieved the highest baseline accuracy of 59.1%. Fine-tuning significantly improved accuracy. For instance, the InternVL3-2B model increased from 58.0% to 64.8%, and the InternVL3-1B improved from 55.7% to 63.5%.
LLM-as-judge metrics also show notable gains after fine-tuning, especially for smaller models. In detail, InternVL3-1B improves in relevance (from 3.44 to 3.70), trustworthiness (from 3.44 to 4.13), clarity (from 4.23 to 4.67), and coherence (from 4.23 to 4.70). Larger models like Qwen2.5-VL-7B and InternVL3-8B show less improvement in these metrics. Overall scores agree with this trend. For example, the InternVL3-1B model’s overall score improves from 3.84 to 4.30, while the InternVL3-2B and InternVL3-8B models’ overall scores remain or decrease slightly after fine-tuning.
This suggests that larger models already possess general reasoning capabilities but still require fine-tuning for domain-specific knowledge. Across all models, relevance scores improved less than other metrics, indicating that baseline models can identify general risk factors but lack the detailed safety knowledge required.
In summary, fine-tuning on the DriveSOTIF dataset enhances both image captioning and VQA. Fine-tuned models generate more detailed, context-aware captions and deliver deeper, scenario-specific reasoning aligned with SOTIF-relevant safety assessments.
@@ V-A Model Deployment and Real-time Performance Evaluation
MLLMs can be deployed in a range of environments, including local devices, cloud platforms, and edge computing systems such as multi-access edge computing (MEC). Popular deployment frameworks include HuggingFace, TensorRT-based solutions [45], and vLLM [46]. In this study, we adopt vLLM for its efficiency and support for advanced scheduling strategies.
@@
Two inference strategies were considered. Sequential inference processes inputs one at a time as they arrive. This provides deterministic, per-sample latencies that are critical for real-time applications. Continuous batching dynamically aggregates requests and schedules token-level processing across partially completed sequences. In this research, we utilize vLLM to implement continuous batching with the goal to maximize throughput while maintaining low latency [46].
@@ V-A1 Sequential Processing
We tested this mode using 111 images from DriveSOTIF, each paired with a prompt for SOTIF risk analysis. Generation time for each image were recorded and presented in Table V-A1. GPU VRAM usage was measured via HuggingFace APIs due to vLLM’s memory pre-allocation behavior.
[b]
@@
    LMDeploy and vLLM are used for InternVL3 deployment.
The peak GPU memory usage and per-image processing time were recorded, and the results are presented in Table V-A1. The analysis shows that quantization reduces GPU memory consumption and improves inference speed. For Qwen2-VL (2B), 4-bit quantization achieves the lowest memory footprint (2.05 GB), while FP8 delivers the fastest inference on both the RTX 3090 and Jetson Orin, with an average latency reduction of 14.7% compared to the full-precision model. A similar trend is observed for Qwen2.5-VL (3B), where 4-bit quantization significantly reduces memory usage (from 7.64 GB to 2.90 GB) and lowers mean inference time from 6.86 s to 5.63 s on Jetson Orin.
While 4-bit quantization is more memory-efficient, it runs slightly slower than FP8 on both platforms (3.56 s vs. 2.93 s on Orin). This is likely due to current limitations in vLLM’s integration with bitsandbytes, which lacks fully optimized GPU kernels for INT4 inference and introduces additional overhead through de-quantization or fallback routines.
Among all evaluated models, InternVL3-1B consistently delivers the most efficient performance, with the lowest latency (0.59 s on RTX 3090, 2.16 s on Jetson Orin) and the smallest memory footprint (1.99 GB). These characteristics make InternVL3-1B particularly well-suited for real-time inference on resource-constrained platforms. In contrast, larger models such as InternVL3-2B demand more memory and exhibit higher latency, limiting their practicality on edge devices.
@@ V-A2 Continuous Processing for high-speed driving
To evaluate continuous inference, we replayed four 30 Hz video streams of high-speed driving scenarios, triggering model inference on every frame. Each frame generates a request to the InternVL3-1B vision-language model, which is deployed on an RTX 3090 with continuous batching enabled  [^4] Please refer to Section V of the supplemental material for results on RTX4090 and dual RTX4090 setups. The system supports multiple requests in parallel, and both queuing and auto-timeout mechanisms are in place: when more requests arrive than the hardware can process immediately, excess requests are placed in a first-in, first-out (FIFO) queue, and any request waiting or processing beyond a set timeout threshold is automatically dropped. We varied the maximum number of in-flight (active) requests from 1 up to 30 (concurrency). For each setting, we recorded the API response time and the number of active requests throughout the experiment (see Figure 3).
@@
We visualized these dynamics by jointly plotting queue wait times and API response times for each concurrency setting (Figure 4). At low concurrency, the system processes requests quickly once they reach the model, but incoming requests rapidly accumulate in the queue, leading to prolonged wait times and, frequently, discarded requests due to timeouts. As concurrency increases, response times become more variable and overall latency rises, but queue wait times decrease since more requests are processed in parallel.
This illustrates a fundamental tradeoff: high concurrency increases per-request processing times but keeps queue wait times low, while low concurrency reduces processing latency but leads to longer queues and more dropped requests under heavy load. In practice, we recommend using low concurrency for critical tasks that require rapid responses and higher concurrency for less time-sensitive workloads. This strategy matches the dual-mode architecture or fast-slow mechanism seen in recent vision language model (VLM) and visual language action (VLA) model research [47, 48]. In these systems, real-time software handles high-speed, routine processing (fast mode), and the VLM or VLA model is triggered when needed to provide semantic reasoning (slow mode).
@@
Figure 4: Tradeoff between API response time and queue waiting time in continuous inference mode.
@@ V-C Results
As shown in Table V-B, in our experiments, fine-tuned MLLMs performed effectively in challenging environments, yet it also highlighted critical areas for enhancement. Notably, in Waterloo’s snowy conditions, the MLLM successfully managed complex scenarios with obscured road markers and temporary obstacles like snowplows and maintenance crews.
@@
However, the model struggled to detect subtle hazards, such as the hidden child in the China experiments. While the model detected unusual patterns associated with the obscured visual cues, it did not accurately capture and interpret the partially or fully obscured child. This highlights a scenario that is challenging even for human drivers. One possible reason is the current capacity of the visual encoders used in the MLLM.
@@ VI-A Implications of Experimental Results
Based on experimental results, the model demonstrates open-world generalization by correctly identifying hazards in unseen scenarios. The model also exhibits causal reasoning by linking the combination of nighttime driving, wet roads, and headlight glare to increased risks. Lastly, it shows contextual understanding by recognizing that pedestrians standing next to snowplows are maintenance workers and providing relevant safety assessments and recommendations.
The computational efficiency reported in Section V.A validates that the above capabilities can be achieved without excessive resource demands. In detail. The fine-tuned InternVL3-1B model requires only 2 GB of GPU memory and can achieve inference times as low as 0.59 seconds per image. This makes it possible to deploy the model onboard vehicles as a lightweight, near-real-time risk assessment module. Doing so removes dependence on external APIs, reduces exposure to privacy and cybersecurity risks, and eliminates the need for continuous internet connectivity.
Pairing a fine-tuned model with a decision-making framework like DriveLLM[20] enables autonomous systems to incorporate SOTIF risks into real-time decision-making. This integration improves safety in complex environments and potentially prevents fatal accidents like the child-hidden-in-the-box scenario, where conventional AD systems may fail to anticipate hidden risks. With iterative dataset expansions and human-in-the-loop annotation, the model can be continuously refined and grounded using SFT and reinforcement learning from human feedback (RLHF) [50]. This adaptive approach enables MLLMs to expand to tasks like complex scenario understanding and ODD assessment.
@@ VI-B Domain Transfer
Results in Table III show that proprietary LLM models can provide accurate and contextually relevant answers to complex, domain-specific SOTIF questions without specialized training. This highlights their effectiveness in zero-shot and one-shot scenarios, where little or no task-specific data is available.
@@ VI-C Hallucination
To address this challenge, we implemented a multi-stage hallucination suppression strategy during the SFT process. Our approach includes (1) a post-processing validation step using an independent LLM to cross-verify generated responses against input images, (2) consistency filtering that removes outputs failing validation checks, and (3) iterative correction of inconsistent annotations. This targeted intervention significantly reduces hallucinated content entering the fine-tuning pipeline, ensuring our dataset maintains grounding in factual, image-based annotations.
However, inference-time hallucination remains a persistent challenge, particularly due to inherent knowledge gaps and biases in pre-trained model backbones. Existing mitigation strategies present distinct trade-offs for real-time applications. Retrieval-augmented generation [52], while effective at grounding responses in factual content, introduces additional inference latency that may be incompatible with the sub-second response requirements of autonomous driving systems. Human feedback mechanisms [53], though widely adopted in industry and demonstrably effective at improving model reliability, face scalability constraints when applied to high-frequency inference scenarios typical in autonomous vehicles. At the moment, one feasible direction is grounding model responses through a human-in-loop semi-auto feedback mechanism. Other methods, like latent space steering [54], can also be explored to reduce hallucination while maintaining real-time performance.
@@ VI-D Future of LLM in SOTIF
As current research on LLMs in SOTIF applications remains limited, we outline several promising directions for both text-based and multimodal LLMs. First, a universal description system for SOTIF scenarios can be developed to provide standardized descriptions. Building on this foundation, custom-trained LLMs can generate JSON-based representations for scenario detection, communication, and risk identification. This standardization would make it easier to share and compare SOTIF scenarios across different research groups and industry applications.
@@ VII Conclusion and Future Work
In this paper, we demonstrate that domain-specific fine-tuning transforms general-purpose MLLMs into specialized models that can analyze, reason about, and react to perception-related SOTIF risks. DriveSOTIF, as the first VQA and image-captioning dataset in the SOTIF domain, enables this adaptation and provides a reproducible framework for evaluation and benchmarking.
Our results can be summarized into three key points. First, fine-tuned MLLMs generate more detailed, context-aware captions and deliver deeper, scenario-specific reasoning aligned with SOTIF-relevant safety assessments. Second, these models demonstrate open-world generalization, causal reasoning, and contextual understanding across regions, weather conditions, and traffic scenarios. Lastly, these models achieve an average inference time of 0.59 seconds per image with only 2 GB of GPU memory, making them suitable for low- to moderate-speed applications like urban navigation, automated parking, and traffic jam assistance.
Deploying these models locally on vehicle hardware enhances driving safety while preserving user privacy and removing dependence on external APIs or cloud connectivity. When integrated into existing decision-making systems, they allow autonomous vehicles to account for SOTIF risk in real time. Crucially, these models can be continuously refined via iterative dataset updates and human feedback. This allows MLLMs to improve over time and expand to more complex tasks like scenario understanding and ODD assessment.
For future work, we plan to improve real-time performance for high-speed safety-critical applications and look into hallucination mitigation during the inference stage. We will also expand the current evaluation to more recent VLM models. At the same time, unifying SOTIF risk analysis and cognition along with action generation using VLA frameworks is a promising direction, given the current trend toward end-to-end autonomous driving systems.